Call for Papers Lake Arrowhead, California, USA,

September 20-23, 2005, http://www.nspw.org

Revised: February 28, 2005

Please note new submission and acceptance dates below.

Background

NSPW is a unique workshop that is devoted to the critical examination of new paradigms in security. We solicit and accept papers on any topic in information security subject to the following caveats:

1. Papers that present a significant shift in thinking about difficult security issues are welcome.

2. Papers that build on a previous shift are also welcome.

3. Contrarian papers that dispute or call into question accepted practice or policy in security are also welcome.

4. We solicit papers that are not technology centric, including those that deal with public policy issues and those that deal with the psychology and sociology of security theory and practice.

5. We discourage papers that represent established or completed works as well as those that substantially overlap other submitted or published papers.

6. We encourage a high level of scholarship on the part of contributors. Authors are expected to be aware of related prior work in their topic area, even if it predates Google. In the course of preparing an NSPW paper, it is far better to read an original source than to cite a text book interpretation of it.

Our program committee particularly looks for new paradigms, innovative approaches to older problems, early thinking on new topics, and controversial issues that might not make it into other conferences but deserve to have their try at shaking and breaking the mold.

Participation in the workshop is limited to authors of accepted papers and conference organizers. Each paper is typically the focus of 45 to 60 minutes of presentation and discussion. Prospective authors are encouraged to submit ideas that might be considered risky in some other forum, and all participants are charged with providing feedback in a constructive manner. The resulting intensive brainstorming has proven to be an excellent medium for furthering the development of these ideas. The proceedings, which are published after the workshop, have consistently benefited from the inclusion of workshop feedback.

Call for Papers

We welcome three categories of submission:

1. Research papers should be of a length commensurate with the novelty of the paradigm and the amount of novel material that the reviewer must assimilate in order to evaluate it.

2. Position papers should be 5 - 10 pages in length and should espouse a well reasoned and carefully documented position on a security related topic that merits challenge and / or discussion.

3. Discussion topic proposals. Discussion topic proposals should include an in-depth description of the topic to be discussed, a convincing argument that the topic will lead to a lively discussion, and supporting materials that can aid in the evaluation of the proposal. The later may include the credentials of the proposed discussants. Discussion topic proposers may want to consider involving conference organizers or previous attendees in their proposals.

Important Dates

Note: these dates have been changed since the original announcement.

• Submission deadline: Monday, 18 April 2005.

• Notification of acceptance: Monday, 13 June, 2005.

Further Information

• Simon Foley, General Chair, University College Cork. s.foley@cs.ucc.ie

• Age Singer, Vice Chair San Diego Supercomputer Center, abe@sdsc.edu

• John McHugh, Program Chair,

SEI/CERT, jmchugh@cert.org

• Bob Blakley, Program co-Chair,

IBM, blakley@us.ibm.com

• Karl Levitt, Local Chair,

UC Davis, levitt@cs.ucdavis.edu

Submission

Submissions must include the following:

1. The submission in PDF format, viewable by Adobe Acrobat reader.

2. A justification for inclusion in NSPW. Specify the category of your submission and describe, in one page or less, why your submission is appropriate for the New Security Paradigms Workshop. A good justification will describe the new paradigm being proposed, explain how it departs from existing theory or practice, and identify those aspects of the status quo it challenges or rejects. The justification is a major factor in determining acceptance.

3. An Attendance Statement specifying how many authors wish to attend the workshop. Accepted papers require the attendance of at least one author. Attendance is limited, and we cannot guarantee space for more than one author.

No submission may have been published elsewhere nor may a similar submission be under consideration for publication or presentation in any other forum during the NSPW review process.

In order to ensure that all papers receive equally strong feedback, all attendees are expected to stay for the entire duration of the workshop. We expect to offer a limited amount of financial aid to those who require it. See http://www.nspw.org for details of the workshop policies and for submission procedures.

Program Committee

Co-Chairs

Bob Blakley IBM Tivoli 13001 Burnet Road, Mail Stop 101/3P-011 Austin, TX 78758

+1 512 286 2440

                                              blakley@us.ibm.com

John McHugh CERT NetSA Carnegie Mellon University Pittsburgh, PA 15313

+1 412 268 7737

                                              jmchugh@cert.org

Program Committee

Gerard Allwein  NRL
Konstantin Beznosov University of British Columbia
Boris Dragovic  CSL, University of Cambridge
Michael Franz  University of CA, Irvine
Carrie Gates  Dalhousie University/CERT NetSA
Steven J. Greenwald Independent Consultant
Vivek Haldar  University of CA, Irvine
Carla Marceau  ATC-NY

Ken Olthoff

Ahmad-Reta Sadeghi University of Bochum
Cristina Serban  AT&T Labs
Tara Whalen  Dalhousie University
Mary Ellen Zurko  IBM

An ACSA Workshop